Suppose 15 manual laborers are waiting for a job. There are six positions open. Now, suppose 10 of the 15 have green cards. If the workers are chosen at random, what is the probability that all of the hired workers have green cards?
Answer by stanbon(75887)

Suppose 15 manual laborers are waiting for a job. There are six positions open. Now, suppose 10 of the 15 have green cards. If the workers are chosen at random, what is the probability that all of the hired workers have green cards?
------
# of ways to succeed:: 10C6 = 10C4 = (10*9*8*7)/(1*2*3*4) = 210
# of possible groups of 6 = 15C6 = 5005
-----
P(6 with green cards) = 210/5005 = 0.0420
=================
Cheers,
Stan H.
=================